OpenStudy (anonymous):

Add. (3x2 + 4x + 7) + (x2 – 6x – 5) 3x2 – 2x + 2 4x2 – 2x + 2 4x2 – 10x + 2 3x2 + 2x – 2

OpenStudy (nurali):

(3x2 + 4x + 7) + (x2 – 6x – 5) 4x2 – 2x + 2

OpenStudy (zehanz):

You can add like terms: 3x²+x² = 4x+(-6x)= 7+(-5)=

OpenStudy (zehanz):

@Nurali: you have to explain, not just give the answer.

OpenStudy (anonymous):

Hey, do you need any help explaining?

OpenStudy (zehanz):

@siegel: I'm afraid we're already beyond that...

OpenStudy (anonymous):

Ok well if you need anything!